OpenStudy (anonymous):

Which of the following is the radical expression of a to the four ninths power ?

OpenStudy (anonymous):

4a9 9a4 fourth root of a to the ninth power ninth root of a to the fourth power

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

why a @vlltimelow

OpenStudy (anonymous):

honestly, I don't know I guessed /:

OpenStudy (aaronq):

a to the four ninths power =\(\huge a^{\frac{4}{9}}\) \(\huge a^{\frac{n}{m}}=\sqrt[m]{x^n}\)

OpenStudy (aaronq):

idk why i changed the a to an x

OpenStudy (anonymous):

The correct answer was ninth root of a to the fourth power